# Guardian: Detecting Robotic Planning and Execution Errors with Vision-Language Models

### v2

@@ III-D Real-Robot, Policy-Driven Data Collection
III-E Dataset Statistics and Evaluation
The resulting datasets, RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, contain balanced success/failure examples across both planning and execution, with fine-grained category labels and reasoning traces. Each dataset is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization. Dataset statistics are shown in Table I.
IV Method: The Guardian Model
Figure 5: Left: Overview of the Guardian model architecture. Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification.
@@ IV-A Problem Formulation
We formulate robot failure detection as a visual question answering problem. For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must not only decide whether the plan is correct but also categorize the type of failure when it occurs:
$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow(B_{\text{plan}},C_{\text{plan}}),$ (1)
where $B_{\text{plan}}\in\{0,1\}$ indicates whether the plan is valid, and $C_{\text{plan}}$ specifies the failure category among the five planning failures constructed in Sec. III-B, shown in Fig. 4.
@@
$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow(B_{\text{exec}},C_{\text{exec}}),$ (2)
where $B_{\text{exec}}\in\{0,1\}$ indicates execution success, and $C_{\text{exec}}$ denotes the execution failure category among the five execution failures constructed in Sec. III-B, shown in Fig. 4.
@@ IV-B Model Architecture
The Guardian model is built upon the state-of-the-art open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining]. As illustrated in Fig. 5 (left), it consists of three key components: a text tokenizer that converts text prompts into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal token sequence to predict the answer.
Rather than concatenating multiple images into a single grid-based image as in AHA [duan2025aha], Guardian processes each image independently through the visual encoder. This design preserves fine-grained spatial details within each image and allows the model to explicitly reason about spatial and temporal changes for more accurate failure detection. Furthermore, unlike SuccessVQA [nvidia2025cosmosreason1physicalcommonsense] and AHA [duan2025aha] that output direct classifications, Guardian leverages an explicit reasoning trace before concluding success or failure.
@@ IV-C Model Training
We fine-tune Guardian on robot failure datasets using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder. Training minimizes cross-entropy loss for next-token prediction.
Although CoT has shown promise to improve performance, it brings additional computation overhead. Inspired by prior work [chen2025trainingstrategiesefficientembodied], we explore three strategies for incorporating CoT into failure detection: (1) Vanilla: a baseline model trained and evaluated to directly predict final answers (A) without CoT; (2) Thinking: the model is trained and inferred with explicit reasoning, always generating CoT before A; (3) Dropout: in training, the model alternates between generating CoT+A and directly predicting A, while at test time, only A is produced.
@@ IV-D Integration into Robotic Manipulation Framework
Guardian can be seamlessly plugged into existing robotic manipulation pipelines as a verification layer without requiring any architectural modification. Without loss of generality, consider a modular robotic manipulation framework. As shown in Fig. 5 (right), Guardian can be inserted at each planning and subtask execution step to detect potential failures. Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery.
@@ V-A Experimental Setup
Evaluation datasets. We evaluate models on our newly constructed benchmarks RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, as well as the existing benchmark RoboFail [liu2023reflect], which is a manually created real-world failure dataset using a UR5 arm with a single camera.
Evaluation metric. We report the mean binary detection accuracy for each dataset (success/failure) in Sec. V-B and discuss multi-class categorization in Sec. V-C.
Implementation details. We trained our models using LoRA (rank 16, effective batch size 16) with the AdamW optimizer (weight decay 0.05), bf16 precision, and a cosine learning rate schedule peaking at $4\times 10^{-5}$. Training was conducted on the RLBench-Fail and BridgeDataV2-Fail sets (unless noted), completing in 5.5 hours on 4×H100 GPUs. During training, we randomly selected one or four views on RLBench-Fail data to mitigate overfitting to the multiple views. The best checkpoint is selected using the validation sets.
TABLE II: Comparison of SOTA failure detection models. Execution and Planning binary accuracies are reported. denotes numbers from the AHA paper. Guardian is fine-tuned from InternVL3-8B on our constructed datasets, boosting the performance of the base model by 23%.
(a) In-Domain datasets.
Model RLBench-Fail BDV2-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.59 0.83 0.75 0.86 GPT4.1 0.63 0.87 0.72 0.86 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.65 0.53 0.58 0.54 Sentinel [agia2024unpackingfailuremodesgenerative] 0.57 - 0.57 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.54 0.60 0.53 0.61 InternVL3-8B 0.59 0.70 0.66 0.71 Guardian-8B-Thinking 0.83 0.87 0.85 0.91
(b) Out-Of-Domain datasets.
Model RoboFail [liu2023reflect] UR5-Fail Exec Plan Exec Plan Large-scale generalist models Qwen3-VL-235B-A22B 0.82 0.70 0.79 0.84 GPT4.1 0.82 0.67 0.79 0.88 Specialized / small models CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 AHA-13B [duan2025aha] 0.64 - - - Sentinel [agia2024unpackingfailuremodesgenerative] 0.79 - 0.74 - Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense] 0.67 0.70 0.58 0.63 InternVL3-8B 0.77 0.53 0.71 0.75 Guardian-8B-Thinking 0.86 0.70 0.77 0.89
@@ V-B Comparison with state of the art
Compared methods. We compare our approach against SOTA general and specialized vision–language baselines. General baselines include small-scale (InternVL3-8B [zhu2025internvl3exploringadvancedtraining]), and large-scale models (Qwen3-VL-235B-A22B-Thinking [qwen3vl2025], GPT4.1 [openai2025gpt41]). Specialized baselines comprise Cosmos-Reason1-7B [nvidia2025cosmosreason1physicalcommonsense], AHA-13B [duan2025aha], Sentinel [agia2024unpackingfailuremodesgenerative], and an MLP baseline using CLIP text-image features [pmlr-v139-radford21a]. We apply test-time CoT to the two large-scale models and Cosmos-Reason, as they are trained to produce reasoning traces and benefit from explicit reasoning at inference. Since AHA-13B is not publicly available, we report its binary accuracy on RoboFail from the original paper [duan2025aha]. Sentinel is limited to execution monitoring and is not evaluated for planning.
Results. Table II reports in-domain (ID) and out-of-domain (OOD) binary accuracies. Table II(a) shows that our Guardian model, fine-tuned from InternVL3-8B on failure datasets, surpasses/matches the larger Qwen3VL and GPT4.1 on in-domain execution and planning sets. In Table II(b) we evaluate models zero-shot on OOD datasets: Guardian attains 0.86/0.70 (Exec/Plan) binary accuracy on RoboFail and 0.77/0.89 on UR5-Fail, outperforming/matching larger general-purpose and specialized VLMs. Sentinel, applied atop Qwen3-VL, underperforms its base VLM, likely due to Sentinel’s rigid CoT prompt. Cosmos-Reason, trained on single-view binary failures, often overlooks fine-grained manipulator/object state changes. AHA, solely tuned to simulation failure data, does not generalize well to the real robot failures of RoboFail.
@@
Training Data RLBench BDV2 Robo UR5 RLBench BDV2 UR5 -Fail -Fail -Fail -Fail 0.65 0.69 0.65 0.73 0.80 0.72 0.69 0.74 0.66 0.87 0.72 0.69 0.85 0.88 0.78 0.83 0.85 0.88 0.74 0.85
Figure 6: Training data size impact on the Execution binary accuracy across benchmarks.
Figure 7: Confusion matrices for Guardian’s failure categorization for plans and executions. Values are percentages; rows denote ground-truth classes and columns the predicted classes.
@@ V-C Failure Data Ablations
In this section, we examine the composition and scale of our data, and discuss the multi-class categorization.
Training data mixture. Table III shows the impact of fine-tuning datasets. Row 1 is the baseline InternVL 3-8B without fine-tuning. Fine-tuning on simulated RLBench-Fail data (row 2) significantly improves the performance on the same dataset, but the gains do not transfer well to the other real-world datasets due to the domain gap. A similar trend is observed when training solely on BridgeDataV2-Fail in row 3. In contrast, combining both datasets during fine-tuning in the fourth row leads to significant gains in domain and improves generalization to the OOD benchmarks. This indicates the importance of generating diverse planning and execution data. Further fine-tuning with the small real-robot UR5-Fail dataset improves performance on UR5-Fail while maintaining comparable accuracy on the other datasets.
Training data size. Fig. 6 presents the scaling behaviors of the training data size on the execution binary accuracies across RLBench-Fail, BridgeDataV2-Fail, RoboFail, and UR5-Fail. We observe a scaling behavior on in-domain and out-of-domain sets when fine-tuning on our procedurally generated data. This suggests that scaling the generated data size further could improve model performance.
Multi-class Categorization. Since our datasets contain fine-grained failure categorization, Guardian supports multi-class failure prediction. The confusion matrices in Fig. 7 show strong overall discrimination across categories. As shown in the Fig. 7(left), most execution categories are identified with high accuracy. The main weakness lies in the VLM’s visual grounding of subtle cues: no gripper close is confused with imprecise grasping/pushing, and wrong object state or placement is often mistaken for success. For planning shown in Fig. 7(right), wrong object state or placement, wrong order, missing subtask, and wrong object manipulated show higher confusion, particularly in long-horizon plans. Improving the model’s commonsense and temporal reasoning would further enhance its ability to capture these nuanced dependencies.
TABLE IV: Image representation impact on the Execution binary accuracy (number of views and image format). The AHA method [duan2025aha] equals row 3 (4-view concatenation).
Views
Image
Format
RLBench-Fail
Execution
1 concat 0.61
separated 0.63
4 concat (AHA [duan2025aha]) 0.72
separated 0.81
Figure 8: Verification with Guardian during online task execution on RLBench. Left: Successful correction of a generated plan. Right: Successful correction of a subtask execution.
V-D Failure Detection Method Ablations
In this section, we compare visual representation choices and reasoning strategies for VLM-based failure detection.
Multi-view images. Table IV investigates the image representation choice on failure detection. We fine-tune and evaluate InternVL3-8B solely on RLBench-Fail execution data. We vary the number of viewpoints (one or four) and the multi-image format, either separated as in Guardian, or concatenated into a single image as in AHA [duan2025aha]. For concatenated inputs, we represent the views as rows, and start/end images as columns. Using multi-view images consistently outperforms their single-view counterpart. In the single-view setting, separating or concatenating two images (start and end) yields comparable performance, as the resolution loss from concatenation is minimal. However, in the four-view setting, separating the images significantly boosts the performance by 9 points over concatenation. Concatenating images not only shifts from natural image distributions but also compresses visual information in limited image resolution.
TABLE V: Comparison of training and test-time strategies with and without CoT (Sec. IV-C). “A” denotes the final answer. We report the average binary accuracy and inference time [seconds/sample] on one H100.

### v3

@@ III-E Dataset Statistics and Evaluation
IV Guardian: A multi-view reasoning VLM for robot failure detection
Figure 4: Left: Overview of the Guardian model architecture. Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification. The failure reasoning trace is reinjected to help with replanning.
@@ IV-A Problem Formulation
We formulate robot failure detection as a visual question answering problem. For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must decide whether the plan is correct or not:
$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow B_{\text{plan}}$ (1)
where $B_{\text{plan}}\in\{0,1\}$ indicates planning success.
@@
$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow B_{\text{exec}}$ (2)
where $B_{\text{exec}}\in\{0,1\}$ indicates execution success.
@@ IV-B Model Architecture
The Guardian model is built upon the open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining]. As shown in Fig. 4 (left), it comprises three components: a text tokenizer that converts text into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal tokens to predict the answer.
Rather than concatenating multiple images into a single grid-based image as in prior work [duan2025aha, ifailsense2026], Guardian processes each image independently through the visual encoder. This design preserves fine-grained spatial details within each image and allows to explicitly reason about spatial and temporal changes for more accurate failure detection. Furthermore, instead of directly outputting classifications [duan2025aha, pmlr-v232-du23b, ifailsense2026], Guardian generates an explicit reasoning trace before concluding success or failure.
@@ IV-C Model Training
We fine-tune Guardian on FailCoT using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder. Training minimizes cross-entropy loss for next-token prediction.
Although CoT has shown promise to improve performance, it brings additional computation overhead. Inspired by prior work [chen2025trainingstrategiesefficientembodied], we explore three strategies for incorporating CoT into failure detection: (1) Vanilla: a baseline model trained and evaluated to directly predict final answers (A) without CoT; (2) Thinking: the model is trained and inferred with explicit reasoning, always generating CoT before A; (3) Dropout: in training, the model alternates between generating CoT+A and directly predicting A, while at test time, only A is produced. Results in Sec. V-C show that adding reasoning traces consistently improves performance. The Thinking strategy performs best but increases inference time, while Dropout offers a better speed-accuracy trade-off.
@@ IV-D Integration into Robotic Manipulation Framework
Guardian can be seamlessly plugged into existing robotic manipulation pipelines as a verification layer without requiring any architectural modification. Without loss of generality, consider a modular robotic manipulation framework. As shown in Fig. 4 (right), Guardian can be inserted at each planning and subtask execution step to detect potential failures. Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery, and use its fine-grained failure reasoning as a hint to better replan.
@@ V-A Experimental Setup
Evaluation datasets. Our main evaluation focuses on three unseen real-world benchmarks: RoboFail [liu2023reflect], UR5-Fail, and RoboVQA [sermanet2024]. RoboFail is a manually curated single-view UR5 failure dataset. UR5-Fail is our constructed multi-view real-robot dataset. RoboVQA (RVQA) is single-view and spans three embodiments: an Everyday Robots mobile manipulator, a human arm, and a human using a grasping tool. [^1] We use the RoboVQA test split restricted to execution success prediction. The original dataset also contains “planning” questions, but these focus on next-action/state prediction rather than plan verification. In ablations, we additionally report results on FailCoT testing splits. We use average classification accuracy as the metric.
Implementation details. We fine-tune models using LoRA (rank 16, effective batch size 16) with AdamW (weight decay 0.05), bf16 precision, and a cosine schedule peaking at $4\times 10^{-5}$. Training is conducted on H100 GPUs using FailCoT unless otherwise specified. For RLBench-Fail, we randomly sample one or four views during training to mitigate view-specific overfitting. The best checkpoint is selected via validation accuracy.
TABLE II: Comparison of failure detection models on unseen real-world benchmarks. Execution and planning accuracies are reported. denotes numbers from the original paper.
Model Trained on RoboFail [liu2023reflect] UR5-Fail RVQA [sermanet2024]
FailCoT Exec Plan Exec Plan Exec
Closed-Source VLM
GPT-4o 0.80 0.67 0.77 0.85 0.79
GPT-4o +Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] 0.80 0.63 0.76 0.62 0.66
Robotic Failure Detection VLMs
RoboFAC-7B [robofac2025] 0.25 0.05 0.54 0.02 0.52
AHA-13B [duan2025aha] 0.64 - - - -
I-Fail-Sense-3B [ifailsense2026] 0.43 0.67 0.47 0.46 0.53
Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.78 0.53 0.59 0.67 0.76
CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 0.52
I-Fail-Sense-3B [ifailsense2026] 0.76 0.52 0.55 0.6 0.58
Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.82 0.70 0.65 0.83 0.77
Guardian-8B 0.86 0.70 0.77 0.89 0.85
@@ V-B Comparison with State of the Art
Compared methods. We compare against GPT-4o [openai2024gpt4ocard] and specialized robotic failure detectors including Cosmos-Reason2-8B [nvidia2026cosmosreason2], AHA-13B [duan2025aha], RoboFAC-7B [robofac2025], I-Fail-Sense-3B [ifailsense2026], Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative], and CLIP+MLP [pmlr-v139-radford21a]. AHA results come from the original paper, as the model is not publicly released. For the other methods, we run their released checkpoints or train models with the released codebase.
Results. Table II reports performance on the test sets. GPT-4o achieves strong performance due to its scale and general reasoning ability. However, applying the Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] self-interrogation prompting degrades accuracy as it constrains the reasoning ability of the original model. Models trained exclusively on simulated failures (RoboFAC [robofac2025] and AHA [duan2025aha]) show limited transfer to real-robot benchmarks. Both rely on simulation-only perturbations, which likely restrict generalization to unseen real-world manipulators and sensor noise. I-Fail-Sense [ifailsense2026] is trained on both simulation and real-world trajectories, but its supervision focuses primarily on semantic misalignment detection rather than structured planning or low-level control failures, which likely limits its performance on the benchmarks. Cosmos-Reason [nvidia2026cosmosreason2] performs competitively, reflecting strong physical reasoning capabilities, but it is optimized for broad embodied reasoning rather than only for failure verification.
Since prior failure detection models are trained on different datasets, we further fine-tune representative open-source models on the same FailCoT dataset to isolate the effects of data and architecture (Table II). We also include a lightweight CLIP+MLP baseline, which performs substantially worse, highlighting the necessity of large vision–language models. Notably, training on FailCoT consistently improves all methods, underscoring the importance of well-curated, cross-environment data with broad failure coverage.
Guardian achieves the strongest overall performance across RoboFail, UR5-Fail, and RoboVQA. Compared to I-Fail-Sense [ifailsense2026], Guardian preserves multi-view spatial structure and produces explicit chain-of-thought reasoning, enabling structured subtask-level verification. Compared to Cosmos-Reason, which is pretrained for embodied reasoning and primarily developed in single-view settings, Guardian leverages an InternVL backbone with explicit multi-view supervision, which likely explains its stronger fine-grained failure detection performance even under identical training data.
@@
Training Data RLBench BDV2 Robo UR5 Robo RLBench BDV2 -Fail -Fail -Fail -Fail -VQA 0.65 0.69 0.65 0.73 0.75 0.82 0.70 0.69 0.72 0.66 0.65 0.86 0.71 0.68 0.77 0.85 0.88 0.78 0.83 0.85
Figure 5: FailCoT training data size impact on the binary accuracy averaged over planning and execution.
@@ V-C Failure Data Ablations
We next analyze how training data composition (simulation and real data), and dataset scale influence cross-environment generalization while keeping the architecture fixed.
Data composition. Table III compares InternVL3-8B without fine-tuning, with single-source training, and with the full FailCoT dataset. Without fine-tuning, performance is moderate. Training only on RLBench-Fail improves simulated results but transfers weakly to real-robot datasets. Training only on BridgeDataV2-Fail (BDV2) improves real-world performance but shows limited simulation transfer. Combining both datasets yields consistent gains across RoboFail, UR5-Fail, and RoboVQA. The same trend holds for other architectures in Table II, emphasizing the importance of cross-environment composition and broad coverage of failure supervision.
Dataset scaling. Fig. 5 demonstrates consistent scaling behavior as the amount of FailCoT data increases. Performance on in-domain and unseen real-world datasets improves steadily with dataset size. This indicates that scaling structured failure generation remains a promising direction for improving cross-environment generalization.
TABLE IV: Comparison of training and test-time strategies with and without CoT (Sec. IV-C). “A” denotes the final answer. We report the binary accuracy and inference time (seconds/sample) on one H100 averaged across all test sets.